INFRARED SPECTROSCOPY OF CH_{2}CL IN SOLID PARAHYDROGEN

\underline{YUKI MIYAMOTO}, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan; and TAKAMASA MOMOSE, Department of Chemistry, The University of British Columbia, Vancouver, B.C., V6T1Z1, Canada.

Infrared spectra of the chloromethyl radical (CH_{2}Cl) isolated in solid parahydrogen has been reported. The radical was produced by in situ UV photolysis of chloroiodomethane (CH_{2}ClI) in solid parahydrogen. Infrared absorption of the ν_{3} (CCl stretch), ν_{2} (CH_{2} scissors), and ν_{1} (CH_{2} symmetric-stretch) modes of CH_{2}Cl were clearly observed. The matrix shift of the ν_{1} band is -6 cm^{−1} relative to the gas phase band origin. Fine spectral features in the infrared absorption spectra reveal that the radical exhibits uni-axial quantized rotation along molecular a-axis in solid parahydrogen. The temporal change of the infrared absorption intensity indicates that the nuclear spin conversion between para (I = 0) and ortho (I = 1) spin modifications of the radical takes place in solid parahydrogen at 3.6 K. The conversion rate was determine to be 0.0046 min^{−1}, which is comparable to the conversion rate between the J = 1, I = 1 and J = 0, I = 2 states of CH_{4} in solid parahydrogen. We did not observe any significant change of the intensity of the CH_{2}Cl radical for a few days in solid parahydrogen, which indicates that there is no tunneling reaction between CH_{2}Cl and H_{2} at 3.6 K.